OpenStudy (anonymous):

Determine the values of b and c so that (3,4) Is the vertex of the given parabola. y=2x^2+bx+C

OpenStudy (amistre64):

vertex x component is given by -b/2a

OpenStudy (amistre64):

\[\frac{-b}{2a}=3\] \[\frac{-b}{2(2)}=3\] \[\frac{-b}{4}=3\] \[-b = 12\] \[b=-12\]

OpenStudy (amistre64):

now we just solve C for y=4 and x=3 y = 2x^2 + bx + C 4 = 2(3^2) -12(3) + C ;

OpenStudy (anonymous):

thanks !